# INDIC QA BENCHMARK: A Multilingual Benchmark to Evaluate Question Answering capability of LLMs for Indic Languages

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable zero-shot and few-shot capabilities in unseen tasks, including context-grounded question answering (QA) in English. However, the evaluation of LLMs’ capabilities in non-English languages for context-based QA is limited by the scarcity of benchmarks in non-English languages. To address this gap, we introduce Indic-QA, the largest publicly available context-grounded question-answering dataset for 11 major Indian languages from two language families. The dataset comprises both extractive and abstractive question-answering tasks and includes existing datasets as well as English QA datasets translated into Indian languages. Additionally, we generate a synthetic dataset using the Gemini model to create question-answer pairs given a passage, which is then manually verified for quality assurance. We evaluate various multilingual Large Language Models and their instruction-fine-tuned variants on the benchmark and observe that their performance is subpar, particularly for low-resource languages. We hope that the release of this dataset will stimulate further research on the question-answering abilities of LLMs for low-resource languages.
@@
Abhishek kumar singh, Rudra Murthy§, Vishwajeet Kumar§ Jaydeep Sen§,Ganesh Ramakrishnan Indian Institute of Technology Bombay, India §IBM Research, India {abhisheksingh,ganesh}@cse.iitb.ac.in, {rmurthyv, vishk024, jaydesen}@in.ibm.com
"In a world deluged by irrelevant information, clarity is power."
– Yuval Noah Harari
1 Introduction
India has the largest global population, with almost 1.43 billion people. However, many of the major languages in India are considered low-resource by the natural language processing (NLP) community. These languages are not as well-represented in NLP compared to English because of the lack of high-quality datasets for pre-training, fine-tuning, and task-specific evaluations. In the field of NLP, Large Language Models (LLMs) have been pre-trained on vast amounts of textual data. Despite their extensive training, these models frequently yield inaccurate results in tasks such as question answering, largely due to their limited contextual comprehension and uncertainties surrounding their parametric knowledge. Researchers have addressed this issue by utilizing systems like Retrieval-Augmented Generation (RAG). RAG retrieves relevant text from a large corpus, whether it’s from pre-training data or real-world data, and uses it as the context for the query. While retrieval is important, the generator component, which finds the exact and correct answer without producing inaccurate outputs, is equally crucial, especially when the answer is not present in the retrieved text.Gao et al. (2023).
Figure 1: One instance from the evaluation benchmark in Hindi languages.
In a Retrieval-Augmented Generation (RAG) system, there are two distinct components: the retriever and the generator. Each component is evaluated separately. The retriever’s performance is assessed through tasks such as passage reranking, utilizing datasets like Thakur et al. (2021), where the retrieval model is trained to rank paragraphs based on their relevance to a given query. On the other hand, the generator’s evaluation focuses on context-grounded question answering. In RAG setups, context is obtained through retrieval, necessitating that the dataset structure includes triples comprising context, question, and answer. The model’s objective is to generate an answer to the query based on the provided context.
@@ 1 Introduction
What distinguishes our benchmark from other existing Multilingual Indic context-grounded question-answering Benchmarks? There are numerous context-grounded question-answering benchmarks available for high-resource languages like English. However, there are very few benchmarks available for Indic languages, and those that do exist often lack domain diversity and are limited in size. To address these gaps, we developed the INDIC QA BENCHMARK, which not only includes a large number of data instances but also spans a wide range of domains, including geography, Indian culture, news, and more. Given the existing domain diversity in English datasets Lee et al. (2019); Bonifacio et al. (2021); Rajpurkar et al. (2016), we decided to translate these datasets into Indic languages. While there are many extractive question-answering datasets, there is a scarcity of abstractive question-answering datasets where the answer may not be explicitly present in the text. To address this, we sampled numerous Wikipedia and Common Crawl pages, focusing on paragraphs rich in cultural nuances and domain diversity. Using these paragraphs, we employed large language models (LLMs) to generate QA pairs, thus creating a comprehensive and culturally diverse benchmark for Indic languages.
In our observations, we have noticed that directly using a base pre-trained model does not initially predict or output answers effectively. Even when it does find an answer, the output is often incorrect and illogical, alongside the ground truth. However, after few-shot prompting, the model produces better-formatted answers and causes the model to start searching for exact answers in single words or phrases from the paragraph, for generation. We summarize the key contributions of this paper as follows:
@@
  • INDIC QA BENCHMARK: a multilingual evaluation benchmark for evaluating Indic Question-Answering/Generative capability of large language models, all these languages are low-resource language and lack a proper benchmark for multi-domain generative tasks.
@@
  • We critically evaluate some of the most esteemed Language Model Models (LLMs) for Indic languages of equivalent size, meticulously comparing their performance on our proposed benchmark to determine their Question-Answering (QA) skills.
@@
Table 1: Indic-QA Dataset Statistics. Indic-QA benchmark is a compilation of existing datasets, English datasets translated to Indian languages, and, synthetic dataset generated using Gemini. The dataset comprises Extractive Question Answering and Generative Question Answering ({*}). As: Assamese, Bn: Bengali, Gu: Gujarati, Hi: Hindi, Kn: Kannada, Ml:Malayalam, Mr: Marathi, Od: Odia, Pa: Punjabi, Ta: Tamil, Te: Telugu. We have translated NQ Open, XORQA, Llama Index, and MS Marco datasets to Hindi. We have translated all the above datasets to the remaining ten Indian languages (underline data instances were already present in the referred language.)
@@ 2 Related Work
In the realm of context-grounded question answering (QA), significant research has been conducted in both English and Indian languages. This task involves presenting a question along with a contextual paragraph to the model, which then extracts the phrase from the paragraph. Various benchmarks Dzendzik et al. (2021); Rajpurkar et al. (2016, 2018) have been established for this task, with encoder-only transformer models proving effective in extracting the span containing the answer from the paragraph. The Indic QA community has demonstrated remarkable performance using models like XLM-RoBERTaConneau et al. (2019) and others, particularly for multilingual Indian languages. They have a rich dataset to showcase their benchmarks, including SQuAD Rajpurkar et al. (2016) for English, along with its translated version in Hindi. Additionally, instead of translation, there are datasets specifically designed for evaluating benchmarks in Hindi, such as the Chaii  [^1] https://www.kaggle.com/competitions/chaii-hindi-and-tamil-question-answering. dataset and IndicQA, which are also discussed in this survey paper Kolhatkar and Verma (2023). Although there are a few benchmarks for Indic Question Answering, they lack extensive domain coverage, which is crucial for evaluating the robustness of models. In contrast, English benchmarks encompass a wide range of domain-specific datasets such as Resources like the llama Index Liu (2022) highlight that selecting the appropriate evaluation dataset is challenging and highly dependent on the specific use case. Academic benchmarks such as BEIR and HotpotQA often fail to generalize across different use cases. For example, parameters that work well on certain data domains (e.g., SEC filings) may not perform as effectively on others (e.g., research papers). This challenge led them to the create a dataset hub specifically designed for evaluating RAG systems, encompassing a wide range of domains including research papers, blockchain articles, and code articles.
@@
3 Benchmarks
The primary focus of this work is on context-based QA, where the answer is entirely or partially found within the given context. The datasets utilized in this study were tailored to facilitate this task, with each instance composed of triples consisting of a context, a question, and an answer. This section provides a detailed description of the methodology used to create or modify the existing dataset for our task.
@@ 3.1 Datasets
    XQuAD: XQuAD (Cross-lingual Question Answering Dataset) Artetxe et al. (2019) serves as a benchmark for assessing the performance of cross-lingual question answering. It comprises 240 paragraphs and 1190 question-answer pairs, extracted from the development set of SQuAD v1.1 Rajpurkar et al. (2016). The dataset includes professional translations of these pairs into ten languages, But we have Used Singh et al. (2024) version of XQuAD because they manually translated to all Indic languages.
@@
    Indic QA Doddapaneni et al. (2022): This dataset is a manually curated cloze-style reading comprehension dataset designed for evaluating question-answering models in 10 Indic languages Since this dataset doesn’t have Gujarati translation we translated it from Hindi to Gujarati and validated the translation as described in the [3.2] section.
@@
    We then translated this dataset from English to Hindi. After applying certain filtering conditions, we translated the Hindi dataset into 10 other Indian languages. The exact steps are detailed in [3.2]. The final dataset now includes the question, the source document, and the corresponding answer, and is available in 11 Indian languages.
@@
    LLama Index Liu (2022): The dataset includes question-answer pairs along with source context, serving as an evaluation tool for the RAG pipeline. We observed that some contexts were insufficient to answer the questions effectively. To address this, we applied the BGE-M3 Chen et al. (2024) algorithm to measure the similarity between the context and the query, using a threshold of 0.43 to determine if a question could be answered adequately based on the context. Post filtering we translated the resulting context, question, and answer triplets into Hindi and Hindi other Indian languages.
@@
    Synthetic Data: This dataset is introduced as part of this study. We employed the Gemini model Team et al. (2023) to generate question-answer pairs based on provided contexts. To achieve this, we sampled a diverse set of Hindi contexts from sources such as Wikipedia, storybooks, Indian news articles, and paragraphs from competitive exams. We then prompted the model with these context paragraphs to generate abstractive question-answer pairs, framing the task as a generative QA task. Subsequently, this dataset was translated into other languages and verified by language experts, the whole workflow process can be found 2.
@@ 3.2 Data Curation Methodology
In light of the approaches discussed previously in Section 2, context-grounded question-answering datasets can generally be categorized into two types: abstractive and extractive. While there are many extractive datasets available in high-resource languages, a few extractive datasets available for Indian languages lack diversity in domains and question types, limiting their utility for benchmarking. Hence, we extended the benchmark suite available in English to these Indian languages by translating. We utilized Indic Transv2  [^2] https://github.com/AI4Bharat/IndicTrans2 Gala et al. (2023) for translation, an open-source transformer-based multilingual NMT model that supports high-quality translations across all the 22 scheduled Indian languages. We segmented the context paragraph into sentences using the Spacy library, translated each sentence, and then recombined them. This approach yielded better translation results, and importantly, the model did not lose context when translating, thus preserving the coherence of the text.
In the list of datasets for benchmarking, some are available only in English (e.g., NQ-open, ORQA, llama index,MS-Marco), while others are available in both English and Hindi (e.g., Hindi SQuAD, CHAII, MLQA, Synthetic data). Additionally, a few datasets (e.g., IndicQA, XSQuAD) are also available in all 10 or 11 languages with verified translations. For all the datasets not found in the respective language, we translated them and applied the filtering methods discussed below.
@@
To assess the quality of our translations, we first translated each dataset from the source language to the target language, then back-translated it from the target language to the source language. We calculated the CHRF and CHRF++ scores between the original and back-translated sentences, applying a threshold on these metrics to filter the instances. Additionally, we manually verified a subset of the filtered data to ensure accuracy. For the translation process, we initially translated the English data directly into Hindi. After filtering the data, we then translated it from Hindi to other Indian languages, rather than directly from English. This approach was based on our observation that the translation quality from Hindi to other Indian languages was superior. The improved quality can be attributed to the linguistic similarities within the same language family, including morphology, syntax, and grammar.

### v2

@@ Abstract
Large Language Models (LLMs) perform well on unseen tasks in English, but their abilities in non-English languages are less explored due to limited benchmarks and training data. To bridge this gap, we introduce the Indic-QA Benchmark, a large dataset for context-grounded question answering in 11 major Indian languages, covering both extractive and abstractive tasks. Evaluations of multilingual LLMs, including instruction fine-tuned versions, revealed weak performance in low-resource languages due to a strong English-language bias in their training data. We also investigated the Translate-Test paradigm,where inputs are translated to English for processing and the results are translated back into the source language for output. This approach outperformed multilingual LLMs, particularly in low-resource settings. By releasing Indic-QA, we aim to promote further research into LLMs’ question-answering capabilities in low-resource languages. This benchmark offers a critical resource to address existing limitations and foster multilingual understanding.  [^1] Source code and Data are available at https://github.com/ayushayush591/IndicQA-Benchmark.
@@
Abhishek kumar singh, Vishwajeet Kumar§, Rudra Murthy§ Jaydeep Sen§, Ashish Mittal§, Ganesh Ramakrishnan Indian Institute of Technology Bombay, India §IBM Research, India {abhisheksingh,ganesh}@cse.iitb.ac.in, {vishk024,rmurthyv,jaydesen,arakeshk}@in.ibm.com
Figure 1: Total tokens available for each Indian language in the Sangraha Data Rahman Khan et al. (2024). In contrast, RefinedWeb Penedo et al. (2023) contains around 5 Trillion tokens in English.
Figure 2: Comparison of LLama 3-8B evaluation results using source language test set vs Translate test set. The results clearly indicate that the Translate-Test paradigm yields better scores for low-resource languages (Punjabi, Odia, Assamese), whereas the source language test set gets better scores for mid-resource language Hindi and Marathi which has high correlation with Hindi.
India, with a population of almost 1.4 billion people, is home to numerous major languages that are considered low-resource by the natural language processing (NLP) community. Despite the growing capabilities of Large Language Models (LLMs) in tasks like context-grounded question answering (CQA) in English, their performance in non-English languages remains underexplored due to a lack of high-quality datasets. To address this gap, we introduce Indic-QA, the largest publicly available context-grounded question-answering dataset for 11 major Indian languages from two language families. This dataset encompasses both extractive and abstractive QA tasks, incorporating existing datasets as well as English QA datasets translated into Indian languages. Additionally, we generate a synthetic dataset using the Gemini model, with manual verification for quality assurance.Our evaluation of various multilingual LLMs and their instruction-fine-tuned variants on the Indic-QA benchmark reveals subpar performance, particularly for low-resource languages. This outcome highlights the English language bias inherent in these models due to predominantly English pre-training data.
We tested the Translate-Test paradigm as an alternative, which translates input from the source language to English, utilizes the LLM’s problem-solving ability in English, and then translates the response back to the source language. Our investigation shows that while multilingual LLMs perform better in mid-resource languages, the Translate-Test paradigm significantly outperforms them in low-resource languages.
@@
What distinguishes our benchmark from other existing Multilingual Indic context-grounded question-answering Benchmarks? There are numerous context-grounded question-answering benchmarks available for high-resource languages like English. However, there are very few benchmarks available for Indic languages[1], and those that do exist often lack domain diversity and are limited in size. To address these gaps, we developed the Indic-QA benchmark. We sampled a variety of Wikipedia and Common Crawl pages, focusing on paragraphs rich in cultural nuances, to create a comprehensive and culturally diverse benchmark for Indic languages. Our findings show that base pre-trained models frequently produce incorrect or illogical answers. However, few-shot prompting notably improves answer quality by guiding the models to extract more precise information from the text. This paper makes the following key contributions:
  1. 1.
@@
    Indic-QA BENCHMARK: We release a multilingual evaluation benchmark for assessing the Indic Question-Answering capabilities of LLMs, focusing on low-resource languages and multi-domain abstractive tasks.
  2. 2.
@@
    Empirical Evaluation: We critically evaluate several esteemed LLMs for Indic languages, comparing their performance on the new benchmark to determine their QA skills.
  3. 3.
    Translate-Test Paradigm: We conduct an empirical study using the Translate-Test approach and direct generation in source languages with multilingual LLMs, demonstrating that the Translate-Test approach offers competitive and often superior performance for low-resource languages.
@@
Table 1: Indic-QA Dataset Statistics. Indic-QA benchmark is a compilation of existing datasets, English datasets translated to Indian languages, and, synthetic dataset generated using Gemini. The dataset comprises Extractive Question Answering and abstractive Question Answering ({*}). As: Assamese, Bn: Bengali, Gu: Gujarati, Hi: Hindi, Kn: Kannada, Ml:Malayalam, Mr: Marathi, Od: Odia, Pa: Punjabi, Ta: Tamil, Te: Telugu. We have translated NQ Open, XORQA, Llama Index, and MS Marco datasets to Hindi. We have translated all the above datasets to the remaining ten Indian languages (underline data instances were already present in the referred language.)
@@ 1 Related work
In the realm of context-grounded question answering (QA), significant research has been conducted in both English and Indian languages. This task involves presenting a question along with a contextual paragraph to the model, which then extracts the phrase from the paragraph. Various benchmarks Dzendzik et al. (2021); Rajpurkar et al. (2016, 2018) have been established for this task, with encoder-only transformer models proving effective in extracting the span containing the answer from the paragraph. The Indic QA community has demonstrated remarkable performance using models like XLM-RoBERTaConneau et al. (2019) and others, particularly for multilingual Indian languages. They have a rich dataset to showcase their benchmarks, including SQuAD Rajpurkar et al. (2016) for English, along with its translated version in Hindi. Additionally, instead of translation, there are datasets specifically designed for evaluating benchmarks in Hindi, such as the Chaii  [^2] https://www.kaggle.com/competitions/chaii-hindi-and-tamil-question-answering. dataset and IndicQA, which are also discussed in this survey paper Kolhatkar and Verma (2023). Although there are a few benchmarks for Indic Question Answering, they lack extensive domain coverage, which is crucial for evaluating the robustness of models. In contrast, English benchmarks encompass a wide range of domain-specific datasets such as Resources like the llama Index Liu (2022) highlight that selecting the appropriate evaluation dataset is challenging and highly dependent on the specific use case. Academic benchmarks such as BEIRThakur et al. (2021) and HotpotQAYang et al. (2018) often fail to generalize across different use cases. For example, parameters that work well on certain data domains (e.g., SEC filings) may not perform as effectively on others (e.g., research papers). This challenge led them to the create a dataset hub specifically designed for evaluating RAG systems, encompassing a wide range of domains including research papers, blockchain articles, and code articles.
@@
Earlier attempts at the Translate-Test approach Etxaniz et al. (2023); Intrator et al. (2024) faced limitations due to less advanced translation systems. However, the emergence of larger bilingual parallel datasets Reid and Artetxe (2022) has allowed researchers to develop robust neural translation models, greatly improving translation performance. To the best of our knowledge, no one has Tried the Translate-Test approach for Indic QA systems.
2 Benchmarks
The primary focus of this work is on context-based QA, where the answer is found within the given context. The datasets utilized in this study were tailored to facilitate this task, with each instance composed of triples consisting of a context, a question, and an answer. This section provides a detailed description of the methodology used to create or modify the existing dataset for our task.
@@ 2.1 Datasets
    XQuAD: XQuAD (Cross-lingual Question Answering Dataset) Artetxe et al. (2019) is a benchmark for evaluating cross-lingual question answering performance. It consists of 240 paragraphs and 1,190 question-answer pairs sourced from the SQuAD v1.1 development set Rajpurkar et al. (2016), with professional translations into ten languages. However, we use the version from Singh et al. (2024), which includes manual translations for all Indic languages.
@@
    Indic QA Doddapaneni et al. (2022): This dataset is a manually curated cloze-style reading comprehension dataset designed for evaluating question-answering models in 10 Indic languages Since this dataset doesn’t have Gujarati translation we translated it from Hindi to Gujarati and validated the translation as described in the [2.2] section.
@@
    We then translated this dataset from English to Hindi. After applying certain filtering conditions, The exact steps are detailed in [2.2]. The final dataset now includes the question, the source document, and the corresponding answer, and is available in 11 Indian languages.
@@
    LLama Index  [^3] https://www.llamaindex.ai/blog/introducing-llama-datasets-aadb9994ad9e Liu (2022): The dataset includes question-answer pairs along with source context, serving as an evaluation tool for the RAG pipeline. We observed that some contexts were insufficient to answer the questions effectively. To address this, we applied the BGE-M3 Chen et al. (2024) algorithm to measure the similarity between the context and the query, using a threshold of 0.43 to determine if a question could be answered adequately based on the context. Post filtering we translated the resulting context, question, and answer triplets into Hindi and Hindi other Indian languages.
@@
    Synthetic Data: This dataset is introduced as part of this study. We employed the Gemini model Team et al. (2023) to generate question-answer pairs based on provided contexts. To achieve this, we sampled a diverse set of Hindi contexts from sources such as Wikipedia, storybooks, Indian news articles, and paragraphs from competitive exams. We then prompted the model with these context paragraphs to generate abstractive question-answer pairs, framing the task as a abstractive QA task. Subsequently, this dataset was translated into other languages and verified by language experts, the whole workflow process can be found [3] [A.1].
Figure 4: Inference in source Language (Top) vs Translate Test Inference (Bottom).
@@ 2.2 Data Curation Methodology
In light of the approaches discussed previously in Section 1, context-grounded question-answering datasets can generally be categorized into two types: abstractive and extractive. Although many extractive datasets exist for high-resource languages, the few available for Indian languages lack diversity in domains and question types, limiting their usefulness for benchmarking. Hence, we extended the benchmark suite available in English to these Indian languages by translating. We utilized IndicTrans2444https://github.com/AI4Bharat/IndicTrans2 Gala et al. (2023) for translation, an open-source transformer-based multilingual NMT model that supports high-quality translations across all the 22 scheduled Indian languages. We segmented the context paragraph into sentences using the Spacy library, translated each sentence, and then recombined them. This approach yielded better translation results, and importantly, the model did not lose context when translating, thus preserving the coherence of the text. In the list of datasets for benchmarking, some are available only in English (e.g., NQ-open, ORQA, llama index,MS-Marco), while others are available in both English and Hindi (e.g., Hindi SQuAD, CHAII, MLQA, Synthetic data). Additionally, a few datasets (e.g., IndicQA, XSQuAD) are also available in all 10 or 11 languages with verified translations. For all the datasets not found in the respective language, we translated them and applied the filtering methods discussed below.
@@
To assess the quality of our translations, we initially translated each dataset from the source language to the target language, followed by back-translation from the target language back to the source language. We then calculated the CHRF scoresPopović (2015) between the original and back-translated sentences, using a threshold of 50 to filter the instances. Additionally, we manually verified a subset of the filtered data to ensure accuracy. For the translation process, we initially translated the English data directly into Hindi. After filtering the data, we then translated it from Hindi to other Indian languages, rather than directly from English. This approach was based on our observation that the translation quality from Hindi to other Indian languages was superior. The improved quality can be attributed to the linguistic similarities within the same language family, including morphology, syntax, and grammar.
Languages Bloom Gemma Llama-3 Llama3.1 Gemma-2
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
As 13.86 16.19 21.70 16.08 19.23 24.14 21.29 22.48 33.01 35.60
Bn 17.84 16.07 21.15 16.69 19.14 23.96 27.61 24.00 38.33 35.11
Gu 13.27 18.59 24.90 25.12 20.27 29.55 24.08 32.11 40.10 44.86
Hi 21.69 19.29 34.37 19.72 41.53 30.02 46.76 26.22 44.46 43.18
Kn 15.63 17.86 24.47 16.31 20.39 25.84 21.31 26.22 35.31 37.98
Ml 19.22 17.96 25.20 17.31 22.80 28.07 31.61 29.49 38.35 41.54
Mr 15.12 18.89 23.96 17.46 36.12 27.54 39.98 31.46 41.89 41.10
Od 11.11 15.09 9.06 15.52 14.20 23.76 12.78 24.37 28.81 36.39
Pa 15.60 20.54 28.43 19.88 21.96 30.54 32.04 29.03 43.37 45.10
Ta 19.96 18.22 22.45 17.42 19.74 27.54 28.04 28.06 39.64 40.29
Te 16.07 17.79 23.93 17.72 13.59 25.37 22.13 27.04 34.97 39.24
Table 2: Performance of both Direct Inference and Translate-Test Inference for various Large Language Models on the Zero-Shot Extractive Indic QA Benchmark. We report the average F1 scores across span-extraction datasets. Instances where Translate-Test outperforms Direct Inference are indicated in bold.
Languages Bloom Openhathi Llama-3 Llama-3.1 Gemma Gemma-2
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
As 5.91 11.61 0.33 3.98 3.97 7.69 4.17 10.13 6.98 7.78 11.32 14.23
Bn 8.21 11.35 0.47 3.58 4.40 7.71 5.10 9.93 6.48 7.22 10.88 13.85
Gu 8.73 12.00 0.12 4.20 6.97 7.98 6.05 10.25 9.06 7.14 12.78 14.64
Hi 8.44 12.14 5.19 4.04 11.17 8.21 11.48 10.66 8.63 7.90 12.33 14.84
Kn 8.19 11.87 0.15 3.76 4.88 7.70 4.71 10.19 7.53 8.04 12.34 14.78
Ml 8.39 11.68 1.01 4.13 6.75 7.66 7.99 9.94 7.14 7.56 13.39 14.27
Mr 8.39 11.38 1.71 3.85 10.08 7.79 9.45 9.93 8.23 7.49 11.98 14.23
Od 6.84 11.58 0.00 3.68 7.01 7.61 7.88 9.93 3.38 6.77 11.70 14.12
Pa 7.91 11.90 0.15 4.18 6.10 8.09 5.81 10.28 9.72 7.84 12.96 14.90
Ta 9.52 12.66 0.47 4.34 3.72 7.95 6.17 10.46 7.74 7.94 13.67 15.38
Te 8.96 12.40 0.19 3.90 3.04 8.10 5.56 10.64 9.34 8.14 13.80 15.19
Table 3: Performance of both Direct Inference and translate test inference of various Large Language Models on Zero-Shot Abstractive INDIC QA BENCHMARK. We report the average numbers (Rouge-L) scores across the languages. Instances where Translate-Test outperforms Direct Inference are indicated in bold.
3 Experiments
We conducted a series of experiments to evaluate the performance of existing LLMs, using NVIDIA A100 GPUs in both 40GB and 80GB variants for our computational needs. Our computational needs signify GPUs both for Translation and evaluation over the models. For inference we utilized VLLM Kwon et al. (2023) which is an open-source library that supports LLM inference efficiently.
We evaluate the following LLMs on our benchmark: OpenHathi [^5] https://huggingface.co/sarvamai/OpenHathi-7B-Hi-v0.1-Base and its instruction-finetuned variant (IFV) known as Airavata Gala et al. (2024), Bloom Le Scao et al. (2022) and its IFV named Bloomz, GemmaTeam et al. (2024), and its instruction fine-tuned variant Gemma-IT. OpenHathi (7B parameter model), was created through continual pre-training on the LLaMA-2 model Touvron et al. (2023). Airavata Gala et al. (2024) (7B parameter model) is an instruction fine-tuned version of OpenHathi. Both OpenHathi and Airavata are specifically trained for Hindi.
Gemma and Gemma-IT (7B parameter models) [^6] https://ai.google.dev/gemma/docs were released by Google. While these models are not specifically trained for Indian languages, they demonstrate multilingual capabilities. Additionally, Aya-8B Aryabumi et al. (2024) is an instruction-tuned model specifically designed for multilingual applications. We also tested Aya-101, another instruction-tuned model based on MT5, which is available in a 13B size. Furthermore, we explored the LLaMA {3, 3.1} and LLaMA {3, 3.1} Instruct models [^7] https://ai.meta.com/blog/meta-llama-3/, which are 8B parameter models and part of the LLaMA family. LLaMA-3 has been trained on data from approximately 30 languages, excluding English. we also used Narvasa2.0 888huggingface.co/Telugu-LLM-Labs/Indic-gemma-7b-finetuned-sft-Navarasa-2.0 and Mistral NeMo  [^9] https://mistral.ai/news/mistral-nemo/ because of their strong multilingual performance. This diverse range of models enables a thorough evaluation of the strengths and performance of our benchmarks across different architectures and training methods.